# femtoPro: Real-time linear and nonlinear optics simulations

### v1

@@ 1 Introduction
Thus, the purpose of the present paper is two-fold: 1) to explain the assumptions underlying the femtoPro software such that users may understand its features and limitations, and 2) to introduce and discuss physical models that can be employed for fast simulations on any platform and for various purposes.
@@
The requirement for a lag-free user experience sets a top priority on the speed of the underlying simulations. This is different from the established simulation tools and protocols listed above. Still, the accuracy needs to be sufficient to provide a realistic simulation of the spatial and temporal propagation of ultrafast laser pulses and their interaction with matter. Some requirements are therefore: We desire to take into account Gaussian beam propagation because finite-focus-size effects are relevant to simulate beam overlap alignment procedures; laser beams may be clipped if they hit the edge of an optical element or an iris; femtosecond pulses are modified by dispersion upon propagation through materials and even more so upon resonant interactions; and second-order processes shall be included to simulate frequency conversion, all in real time as defined above.
@@
The underlying simulation concepts are in principle known from the vast existing scientific literature, in particular from textbooks on general optics [18, 19, 20, 21, 22, 23] and ultrafast spectroscopy [24, 25, 26, 27, 28, 29, 30, 31, 32]. However, the real-time simulation requirement necessitates a careful balance between speed and accuracy. In the present work, we derive simplified expressions that facilitate real-time optics simulations.
@@
We start by defining our representation of the electric fields of laser pulses (Section 2). Then we discuss the spatial modulations of laser beams resulting from geometric effects of optical elements and multi-beam interference (Section 3), followed by spatial and spectral–temporal modulations arising from linear and nonlinear light–matter interaction (Section 4) and the graphical representation of laser beams (Section 5). Exemplary simulation results are shown in Section 6 before we conclude in Section 7.
2 Laser pulses
The representation of electric fields of laser pulses is treated in many comprehensive books on general optics [18, 19, 20, 21, 22, 23] and ultrafast time-resolved spectroscopy [24, 25, 26, 27, 28, 29, 30, 31, 32]. Here we seek a representation that is suitable for efficient calculations in computational models. Since we need to describe both spatial and spectral–temporal properties, we have to separate the various aspects in a convenient manner. We define and derive all relevant expressions in great detail in Section LABEL:SISecElectricField of Supplementary Information (SI) and only quote the main results here.
@@ 2 Laser pulses
according to Eq. (LABEL:SIEqEPropXYZT) in the SI, where the “$+$” superscript in $E_{\rm{prop}}^{+}$ indicates that this is a complex-valued representation resulting from the positive-frequency part of the full field only and the subscript “prop” indicates that the pulse has propagated through space along the $\hat{\mathbf{z}}$ direction requiring the propagation time $T$. The beam radius at each $z$ position is given by $w(z)$ (with details on the evolution of $w(z)$ as a function of $z$ provided in Section 5) and lateral translations of the beam along $\hat{\mathbf{x}}$ and $\hat{\mathbf{y}}$ by $x_{0}$ and $y_{0}$, respectively. Further, $\mathbf{k}$ denotes the wave vector, $\omega_{0}$ the center frequency, $\delta t$ the sampling step size in time domain, $S$ a scaling factor proportional to pulse energy, and $\tilde{E}_{t}(j)$ the numerical array representation of the complex-valued temporal envelope (including amplitude and phase terms) for individual sampling points that are indexed by $j$.
@@ 3.1 Finite aperture
Apertures lead to the clipping of laser beams. Recently, analytical propagation formulas have been derived for truncated Gaussian beams [33]. Here we ignore diffraction for reasons of computational speed. If we removed the Gaussian beam properties (2) at the same time, that would correspond to a transition from wave optics to the limit of geometrical optics. The latter is not sufficient, however, if we want to describe interference phenomena and allow for the quantitative treatment of spatial beam overlap effects in nonlinear optical phenomena. Thus, in the present work we suggest an intermediate regime where we treat some phenomena in the limit of geometrical optics and some phenomena using wave optics.
@@
Concerning apertures, we select a treatment by geometrical optics. This has the advantage that we do not have to follow various diffracted beamlets that might propagate in various different directions, in particular after interaction with subsequent optical elements. Instead, we can continue to describe the transmitted beam as a single entity with one limited set of parameters. While this is a limitation that removes some phenomena from being treated with the model, it strongly reduces complexity and facilitates real-time simulations.
@@
Despite this approximation, we seek a representation that is as faithful to reality as possible under this approximation. Thus, a laser beam emerging from any open aperture should not have a cross section that extends beyond the hard limits of the aperture. Such an aperture could for example be an iris that is opened or closed by users either to clip laser beams on purpose or to act as alignment tools. The laser beam then would be centered onto at least two subsequent irises, and after the alignment procedure is completed, the irises would be opened to let the full beam pass. In that case, then, the final beam propagation would again not be limited by diffraction for the “real” experiment, justifying the limit of geometrical optics, while the clipping effects will be captured at least qualitatively or semi-quantitatively (to be discussed below) during the simulation of alignment procedures. As a second possibility, finite-aperture effects become relevant when a laser beam is larger than the optical element it hits, or when it hits the element at its edge. This might happen for misaligned mirrors or lenses that are hit by the laser at their edges instead of in the center. While such situations should be simulated at least qualitatively in a real-time optics education scenario, they should typically not arise in a final, correctly set up experiment. Thus, again, it seems justified to employ a simplified geometrical optics treatment.
The approximation of geometrical optics entails that our model cannot be used to simulate diffraction gratings. This is a current limitation (see Section LABEL:SISecFeaturesAndLimitations in the SI for an overview of features and limitations).
In order to ensure that the Gaussian beam that is transmitted through an aperture does not have a cross section that extends beyond the aperture limits, one has to evaluate the geometrical common cross section of the laser beam and the aperture. At normal incidence, one requires the common area of two circles formed by the aperture and the laser cross section, in which one then has to fit a third circle that does not extend beyond the common limits. For nonlinear signal generation from at least two separate beams, however, one would already have to take into account at least three circles, i.e., one for the aperture and two for the incident beams. The problem of the common intersection of three circles is already quite involved with an analytical solution reported only in 2006 [34]. Anyway, the laser beam will, in general, not hit the aperture at normal incidence. In that case, one would have to consider the common overlap of three ellipses that is even more complicated than that of three circles because of the many types of overlap cases that have to be distinguished [35, 36]. Since we ignore astigmatism, however, it is enough to consider a suitably adapted scenario with circular symmetry.
We have adopted the “mixed” solution that evaluates overlaps between several beams, such as in interference evaluation or nonlinear signal generation, on the basis of overlapping Gaussian functions; as a second component of our approach, we then treat the transmission through a hard aperture in a purely geometrical fashion to ensure that “no” laser beam profile extends beyond the open aperture. The term “no” is set in quotation marks here because the visual appearance of the laser beam, when rendered as a three-dimensional object up to its beam radius, will indeed not be larger than the transmission aperture. On the other hand, since we treat the beam as a Gaussian, this still allows its radial distribution to extend further outwards than the aperture opening. This is a consequence of the hybrid approach, adopted in this work, between geometrical optics and wave optics when diffraction is ignored. Thus, we take the beam radius as the geometric limit. When we then choose to visualize laser beams via plotting their beam radius, an intuitively consistent behavior emerges if the beam waist is taken to represent the “edge.”
@@
For treating the transmission through optics, we have to consider finite incidence angles. For example, mirrors are routinely employed at $\approx 45^{\circ}$ incidence angle to deflect beams, and anyway we cannot avoid that the beam hits a general optical element (GOE) at an arbitrary angle, so the model needs to deal with all situations. Thus we may not assume the “paraxial” limit, and we employ vector calculus for an analysis independent of any particular chosen coordinate system.
The relevant geometry is illustrated in Fig. 1. The incident laser beam (from the left, potentially arriving at an additional angle elevated out of the drawing plane) is given by a point of origin in world coordinates, $\mathbf{r}_{\text{in}}$, in our evaluation logic usually its starting point at the previous GOE; a unit directional vector $\hat{\mathbf{k}}_{\text{in}}=\mathbf{k}_{\text{in}}/\left\lvert\mathbf{k}_{\text{in}}\right\rvert$, where $\mathbf{k}_{\text{in}}$ is the incident wave vector; and a beam radius $w_{\text{in}}$ at the point of intersection with the current GOE. The current GOE is defined by its aperture center in world coordinates, $\mathbf{o}$, a normal vector describing the orientation of the “active” aperture plane, $\hat{\mathbf{n}}$, and an aperture radius $a$.
Figure 1: Aperture geometry. The input beam (red) propagates along a central “ray” (solid) with lateral width determined by its beam radius (dashed–dotted) until it intersects at an incidence angle $\alpha$ with the central active plane of a general optical element (GOE) that is assumed, without loss of generality, to be normal to the “optical axis” (dashed black). The output beam (green) is likewise defined by a central ray (solid) and radius (dashed–dotted). All relevant scalar and vectorial quantities are marked in blue as defined in the text.
We first determine the propagation vector, $\mathbf{L}$, that extends from the origin of the laser beam at the previous GOE to the intersection with the active plane of the current GOE [37],
$\mathbf{L}=\frac{(\mathbf{o}-\mathbf{r}_{\text{in}})\cdot\hat{\mathbf{n}}}{\hat{\mathbf{k}}_{\text{in}}\cdot\hat{\mathbf{n}}}\,\hat{\mathbf{k}}_{\text{in}}\quad\text{if }\hat{\mathbf{k}}_{\text{in}}\cdot\hat{\mathbf{n}}\neq 0.$ (3)
If $\hat{\mathbf{k}}_{\text{in}}\cdot\hat{\mathbf{n}}=0$, the laser beam propagates parallel to the GOE’s active plane, and we have to distinguish further cases. If, in that case, the numerator in Eq. (3) is not equal to zero, the laser beam is parallel to but outside of the surface, so that there is no intersection. If the numerator and the denominator in Eq. (3) are equal to zero, then the laser beam propagates within the surface, i.e., the GOE is hit exactly from the side. We then define the propagation vector such that it propagates the laser beam from its origin to the location closest to the center of the GOE,
$\mathbf{L}=\mathbf{o}-\mathbf{r}_{\text{in}}-\text{proj}_{\hat{\mathbf{k}}_{\text{in}}}(\mathbf{o}-\mathbf{r}_{\text{in}}),$ (4)
with the vector projection
$\text{proj}_{\mathbf{b}}\mathbf{a}=\left(\mathbf{a}\cdot\frac{\mathbf{b}}{\left\lvert\mathbf{b}\right\rvert}\right)\frac{\mathbf{b}}{\left\lvert\mathbf{b}\right\rvert}$ (5)
of vector $\mathbf{a}$ onto vector $\mathbf{b}$.
The propagation length $L=\left\lvert\mathbf{L}\right\rvert$ enters Eq. (LABEL:SIEqPropagationTime) in the SI for the calculation of the propagation time $T$. From the previous origin of the laser and the propagation vector we obtain the “hit” point of intersection with the current GOE, in world coordinates, as
$\mathbf{h}=\mathbf{r}_{\text{in}}+\mathbf{L},$ (6)
and as
$\mathbf{d}=\mathbf{h}-\mathbf{o}$ (7)
in a GOE-local coordinate system with the origin at the center of the GOE, where we further define the scalar intersection distance $d=\left\lvert\mathbf{d}\right\rvert$ from the GOE center. The vectorial quantity $\mathbf{d}$ enters the calculation of directional change by a GOE with focal length not equal to zero, as will be derived below. In the present context, we project the incident beam radius, $w_{\text{in}}$, along direction $\mathbf{d}$ to obtain the projected beam radius, $b_{\text{in}}$. For this, we observe in Fig. 1 that $\cos\alpha=w_{\text{in}}/b_{\text{in}}$ but also that $\cos\alpha=\hat{\mathbf{n}}\cdot\hat{\mathbf{k}}_{\text{in}}$. Eliminating $\alpha$ provides
$b_{\text{in}}=\frac{w_{\text{in}}}{\left\lvert\hat{\mathbf{n}}\cdot\hat{\mathbf{k}}_{\text{in}}\right\rvert},$ (8)
where we introduced the absolute magnitude operator such that it works also if the surface normal $\hat{\mathbf{n}}$ is inverted. In the case of $\hat{\mathbf{n}}\cdot\hat{\mathbf{k}}_{\text{in}}=0$, as noted above, the beam travels within the plane, and the projected radius is infinite as it should be. This leads to the further treatment as case 2 in the list below.
Now we have obtained the three quantities, aperture radius $a$, intersection distance $d$, and projected beam radius $b_{\text{in}}$, that are required for the following case distinction based on their quantitative relation:
  1. 1.
    $a\geq d+b_{\text{in}}$: The laser beam is completely within the aperture;
  2. 2.
    $b_{\text{in}}>d+a$: The laser beam overfills completely the aperture;
  3. 3.
    $d>b_{\text{in}}+a$: The laser beam is completely outside the aperture;
  4. 4.
    Else: The laser beam is partially within the aperture.
Let us evaluate, for each case, the projected radius of the new beam after the aperture, $b_{\text{out}}$, along direction $\mathbf{d}$; the origin of the new beam after the aperture in world coordinates, $\mathbf{r}_{\text{out}}$; and the direction of the new beam after the aperture, $\hat{\mathbf{k}}_{\text{out}}$. Finally, we obtain the new beam radius perpendicular to the new direction of propagation, $w_{\text{out}}$, that is given in all cases, analogous to Eq. (8), by
$w_{\text{out}}=b_{\text{out}}\left\lvert\hat{\mathbf{n}}\cdot\hat{\mathbf{k}}_{\text{out}}\right\rvert.$ (9)
Note that in such a treatment, we consider the beam radius as projected along direction $\mathbf{d}$ only, and we ignore the shape and extension in the orthogonal direction. This is an approximation that works well for meridional rays but less so for skew rays. Nevertheless, we follow this approach because of our choice to ignore astigmatism for computational efficiency [14], and thus one may define just one particular projection coordinate (not two) to represent a beam radius and shape. Using the meridional-ray approximation leads to the desired effect that no transmitted beam is larger than any circular open aperture.
Case 1 ($a\geq d+b_{\text{in}}$) in the list above corresponds to the “normal” alignment goal of having an unclipped beam with an open iris aperture (or a fully reflected beam off a mirror). Then the starting point of the new beam is equal to the intersection point of the incident beam,
$\mathbf{r}_{\text{out}}=\mathbf{h},$ (10)
the new beam direction is given by [38]
$\hat{\mathbf{k}}_{\text{out}}=\frac{\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{d}}{f}}{\left\lvert\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{d}}{f}\right\rvert},$ (11)
and the projected radius is simply
$b_{\text{out}}=b_{\text{in}}.$ (12)
In case 2 ($b_{\text{in}}>d+a$), the clipped beam passes, symmetrically distributed, exactly through the center of the GOE, independently of the local intersection coordinate $\mathbf{d}$ that was calculated for the center of the unclipped beam. Note that in this geometric treatment, as motivated at the beginning of this section, we do not take into account the Gaussian variation of the intensity within the lateral cross section, and thus we arrive at the statement of the symmetric distribution around the GOE center $\mathbf{o}$. Thus, this case corresponds to a “center ray” in geometrical optics that is unaffected by any focal length, and hence we obtain
$\mathbf{r}_{\text{out}}=\mathbf{o}$ (13)
for the origin,
$\hat{\mathbf{k}}_{\text{out}}=\hat{\mathbf{k}}_{\text{in}}$ (14)
for the direction, and
$b_{\text{out}}=a$ (15)
for the projected radius of the new beam.
Case 3 ($d>b_{\text{in}}+a$) is simple to treat because then no beam is transferred.
Finally, in case 4 (“else”), the beam is clipped partially. From Fig. 1, one finds $2b_{\text{out}}=a-d+b_{\text{in}}$ such that the projected radius is
$b_{\text{out}}=\frac{1}{2}(a-d+b_{\text{in}}).$ (16)
If the beam shall indeed arise from the common overlap between original projected radius and aperture radius, we have to laterally “move” the new beam’s origin such that it is in the center of the new beam. From Fig. 1, one obtains $p=\left\lvert\mathbf{p}\right\rvert=d-b_{\text{in}}+b_{\text{out}}=\frac{1}{2}(a+d-b_{\text{in}})$ and, as a GOE-local vector,
$\mathbf{p}=p\frac{\mathbf{d}}{d}=\frac{1}{2}(d+a-b_{\text{in}})\frac{\mathbf{d}}{d}$ (17)
such that
$\mathbf{r}_{\text{out}}=\mathbf{o}+\mathbf{p}$ (18)
and
$\hat{\mathbf{k}}_{\text{out}}=\frac{\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{p}}{f}}{\left\lvert\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{p}}{f}\right\rvert}.$ (19)
Note that now the effective “distance” entering in Eq. (19) is $\mathbf{p}$ rather than $\mathbf{d}$ as in Eq. (11). We illustrate the case of partial beam clipping in Fig. 2, where we vary the lateral shift of a beam incident on a GOE with a finite aperture while displaying the waist and the angle of an output beam. The radii of the input beam and the aperture are set to 3 mm. To invoke a change of the angle of the output beam, we set the focal length of the GOE to $f=37.5$ cm. We display two subcases. In case of a small angle of incidence (5.71), the output beam waist decreases linearly when shifting the input beam away from the center of the aperture [Fig. 2(a)]. The output beam angle also changes linearly. In case of a steeper angle of incidence (45) as shown in Fig. 2(b), the radius and angle of the outgoing beam remain constant in a region around the center of the aperture. When the incident beam is further laterally shifted, clipping again results in a change of radius and angle of the outgoing beam.
Figure 2: Output beam from a finite aperture. The lateral shift of an input beam relative to the optical axis is varied under a given incidence angle. The beam waist of the input beam and the aperture radius are both chosen to be 3 mm while the focal length of the GOE is set to 37.5 cm. The beam waist and the angle of the output beam is shown for an incidence angle of (a) 5.71 and (b) 45. The incidence angles are indicated as horizontal dashed lines.
Finally, we need to consider the reduction in pulse energy due to beam clipping caused by the finite aperture. Above, we made the approximation of uniform spatial profiles for generating the geometric parameters of the new beam. Let us first consider what would happen at the same level of approximation for the transmitted pulse energy (which, as we will see, is not what we implemented). In this simplest approximation of a “flat-top” beam, we could relate the output energy, $W_{\text{out}}$, to the incident pulse energy, $W_{\text{in}}$, by considering the ratio of the projected cross-sectional areas at the aperture,
$W_{\text{out}}=\frac{\pi b_{\text{out}}^{2}}{\pi b_{\text{in}}^{2}}W_{\text{in}}.$ (20)
Note that one should not use the orthogonal beam radii $w_{\text{in}}$ and $w_{\text{out}}$ in this context because they are different already if the output beam direction $\hat{\mathbf{k}}_{\text{out}}$ is different from the incidence beam’s direction $\hat{\mathbf{k}}_{\text{in}}$ due to an off-axis focusing element with focal length $f\neq 0$, i.e., the pulse energy does not change but only the intensity changes due to the changed area.
While Eq. (20) describes the situation for a uniform spatial intensity profile, this does not correspond to the properties of a Gaussian beam in which the intensity is higher in the center than at the edge and which is what we introduced in Eq. (1) and Eq. (2). The discrepancy would be noticeable and is relevant if the simulation model is used, for example, for didactic training in the context of beam alignment onto an aperture. In those cases, the goal is to align the position of a laser beam with the center of a given aperture (or vice versa, center an aperture on a given laser beam) by monitoring the energy changes detected with a power meter after the aperture when moving the beam (or aperture) laterally. Assume now that the laser overfills the aperture (case 2 in the list above). In that case, the transmitted cross-sectional area would not change, and neither would the transmitted pulse energy according to simplistic Eq. (20). Thus, we seek a more accurate calculation of the transmitted energy.
More accurately, we might alternatively consider, instead of Eq. (20) from above, equation (14) from reference [39] that derives, for normal incidence at which $b_{\text{in}}=w$, the following relation:
$W_{\text{out}}=W_{\text{in}}\exp\left(-\frac{2d^{2}}{w^{2}}\right)\times\sum_{k=0}^{\infty}\left\{\frac{2^{k}d^{2k}}{w^{2k}k!}\left[1-\exp\left(-\frac{2a^{2}}{w^{2}}\right)\sum_{i=0}^{k}\frac{2^{i}a^{2i}}{w^{2i}i!}\right]\right\}.$ (21)
It is discussed in reference [39] that for central beams, i.e., $d=0$, one obtains the more well-known relation
$W_{\text{out}}(d=0)=W_{\text{in}}\left[1-\exp\left(-\frac{2a^{2}}{w^{2}}\right)\right],$ (22)
but this, on its own, is not a useful limit for us either because we want to deal with the particular situation of laterally displaced beams.
Thus we evaluate the transmitted energy numerically using precalculated lookup tables as derived in Section LABEL:SISecGaussianTransmissionThroughCircularAperture in the SI. This allows us to assign transmission values fast for any given combination of beam radius $w$, aperture radius $a$, and displacement $d$. Note that we do not change the pulse energy directly, but rather the pulse-energy scaling factor $S$ from Eq. (LABEL:SIEqEnergyIsScaleTimesNormSquared). The scaling factor as a function of the aperture radius is displayed in Fig. 3(a), where the behavior of Gaussian beams is evident by an $S$ less than one in the case of $w=a$. In Fig. 3(b), we laterally shift the input beam relative to an aperture with size $a=3w$. Note that we capture the beams geometrically up to their beam radius, i.e., we treat Gaussian beams, which theoretically go to infinity, as hard-cut structures. This means that if the beam is shifted away from the edge of the aperture by a distance larger than $w$, it is treated as no longer hitting the aperture, leading to zero energy which in turn results in the small “step” visible in Fig. 3(b) at a lateral shift of $4w$. This provides a quantitative estimate for the error at the chosen level of approximation of our model.
Figure 3: Pulse-energy scaling factor S of a Gaussian beam incident on an aperture. (a) Scaling factor as a function of the aperture size. (b) Scaling factor as a function of the lateral shift relative to an aperture with a size of $3w$.
3.2 N-beam interference

### v2

@@ 1 Introduction
Thus, the purpose of the present paper is two-fold: 1) to explain the assumptions underlying the femtoPro software such that users may understand its features and limitations, and 2) to introduce and discuss physical models that can be employed for fast optics simulations on any platform and for various purposes.
@@
The requirement for a lag-free user experience sets a top priority on the speed of the simulations. This is different from the established simulation tools and protocols listed above. Still, the accuracy needs to be sufficient to provide a realistic simulation of the spatial and temporal propagation of ultrafast laser pulses and their interaction with matter. Some requirements are therefore: We desire to take into account Gaussian beam propagation because finite-focus-size effects are relevant to simulate beam overlap alignment procedures; laser beams may be clipped if they hit the edge of an optical element or an iris; femtosecond pulses are modified by dispersion upon propagation through materials and upon resonant interactions; and second-order processes shall be included to simulate frequency conversion, all in real time as defined above.
@@
The underlying simulation concepts are in principle known from the vast existing scientific literature, in particular from textbooks on general optics [18, 19, 20, 21, 22, 23] and ultrafast spectroscopy [24, 25, 26, 27, 28, 29, 30, 31, 32]. However, the real-time simulation requirement necessitates a careful balance between speed and accuracy. In another manuscript, we describe the concept, implementation, and evaluation of a dynamic graph-based algorithm that is used to find a self-consistent solution of multiple propagating and interfering laser beams [33]. The dynamic graph model maps the formal network between beam segments of laser pulses propagating freely in space as “edges,” and optical elements (such as mirrors or lenses) that modify the laser pulse properties as “nodes,” and then the algorithm is set up to minimize the number of evaluation steps. In the present work, we focus on the physical model underlying the individual calculation steps of that algorithm and derive simplified mathematical expressions that facilitate real-time optics simulations.
@@
We start by defining our representation of the electric fields of laser pulses including the spatial modulations of laser beams resulting from geometric effects of optical elements (Section 2). Then we discuss multi-beam interference (Section 3), followed by spatial and spectral–temporal modulations arising from linear and nonlinear light–matter interaction (Section 4) and the graphical representation of laser beams (Section 5). Exemplary simulation results are shown in Section 6 before we conclude in Section 7.
2 Laser pulse propagation
We define and derive all relevant expressions for the description of laser pulses in great detail in Section LABEL:SISecElectricField of Supplementary Information (SI) and only quote the main results here. Indeed, we highly recommend reading the entire Section LABEL:SISecLaserPulsePropagationAndModificationByOpticalElements of the SI before continuing with the main article here because all the relevant quantities and essential transformations are set up in the SI.
@@ 2 Laser pulse propagation
according to Eq. (LABEL:SIEqEPropXYZT) in the SI, where the “$+$” superscript in $E_{\rm{prop}}^{+}$ indicates that this is a complex-valued representation resulting from the positive-frequency part of the full field only and the subscript “prop” indicates that the pulse has propagated through space along the $\hat{\mathbf{z}}$ direction requiring the propagation time $T$. The beam radius at each $z$ position is given by $w(z)$ [Eq. (LABEL:SIEqBeamRadiusEvolution)] and lateral translations of the beam along $\hat{\mathbf{x}}$ and $\hat{\mathbf{y}}$ by $x_{0}$ and $y_{0}$, respectively. Further, $\mathbf{k}$ denotes the wave vector, $\omega_{0}$ the center frequency, $\delta t$ the sampling step size in time domain, $S$ a scaling factor proportional to pulse energy, and $\tilde{E}_{t}(j)$ the numerical array representation of the complex-valued temporal envelope (including amplitude and phase terms) for individual sampling points that are indexed by $j$.
@@
We now discuss how to describe the propagation of laser beams in the presence of finite apertures. Apertures lead to the clipping of laser beams. Recently, analytical propagation formulas have been derived for truncated Gaussian beams [34]. Here we ignore diffraction for reasons of computational speed. If we removed the Gaussian beam properties (2) at the same time, that would correspond to a transition from wave optics to the limit of geometrical optics. The latter is not sufficient, however, if we want to describe interference phenomena and allow for the quantitative treatment of spatial beam overlap effects in nonlinear optical phenomena. Thus, in the present work we suggest an intermediate regime where we treat some phenomena in the limit of geometrical optics and some phenomena using wave optics.
@@
Concerning apertures, we select a treatment by geometrical optics as discussed in detail in Section LABEL:SISecFiniteAperture of the SI. This has the advantage that we do not have to follow various diffracted beamlets that might propagate in various different directions, in particular after interaction with subsequent optical elements. Instead, we can continue to describe the transmitted beam as a single entity with one limited set of parameters. The approximation of geometrical optics entails that our model cannot be used to simulate diffraction gratings. On the other hand, this strongly reduces complexity and facilitates real-time simulations.
@@
We seek a representation that is as faithful to reality as possible under the given model assumptions. Thus, a laser beam emerging from any open aperture should not have a cross section that extends beyond the hard limits of the aperture. Such an aperture could for example be an iris that is opened or closed by users either to clip laser beams on purpose or to act as alignment tools. The laser beam then would be centered onto at least two subsequent irises, and after the alignment procedure is completed, the irises would be opened to let the full beam pass. In that case, then, the final beam propagation would again not be limited by diffraction for the “real” experiment, justifying the limit of geometrical optics, while the clipping effects will be captured at least qualitatively or semi-quantitatively during the simulation of alignment procedures. As a second possibility, finite-aperture effects become relevant when a laser beam is larger than the optical element it hits, or when it hits the element at its edge. This might happen for misaligned mirrors or lenses that are hit by the laser at their edges instead of in the center. While such situations should be simulated at least qualitatively in a real-time optics education scenario, they should typically not arise in a final, correctly set up experiment. Thus, again, it seems justified to employ a simplified geometrical optics treatment.
@@
For treating the transmission through optics, we have to consider finite incidence angles. For example, mirrors are routinely employed at $\approx 45^{\circ}$ incidence angle to deflect beams, and anyway we cannot avoid that the beam hits a general optical element (GOE) at an arbitrary angle, so the model needs to deal with all situations. Thus we may not assume the paraxial limit, and we employ vector calculus for an analysis independent of any particular chosen coordinate system.
In the case of an unclipped beam transmitted through a GOE with focal length $f$, a beam with normalized incident wave vector $\hat{\mathbf{k}}_{\text{in}}$ that hits the GOE at a displacement $d$ with respect to the GOE center will exit the element in a new direction given by the normalized outgoing wave vector $\hat{\mathbf{k}}_{\text{in}}$, where these quantities are related by [35]
$\hat{\mathbf{k}}_{\text{out}}=\frac{\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{d}}{f}}{\left\lvert\hat{\mathbf{k}}_{\text{in}}-\frac{\mathbf{d}}{f}\right\rvert}$ (3)
as derived in Eq. (LABEL:SIEqDirectionChangeGOELaserCompletelyWithinAperture). The beam radius projected onto the plane of the GOE remains the same because there is no clipping.
In the case of partial beam clipping, the quantity $\mathbf{d}$ in Eq. (3) needs to be replaced by an effective quantity $\mathbf{p}$ as in Eq. (LABEL:SIEqDefinitionP) that points from the GOE center to the center of the emerging clipped beam, leading to a modified direction of the outgoing beam according to Eq. (LABEL:SIEqDirectionChangeGOELaserPartiallyWithinAperture).
In the case that the beam is fully clipped, the center of the transmitted beam will be at the center of the GOE even if the incident beam was laterally displaced. This is because we consider only circularly symmetric apertures such that emerging beams that are fully defined by the aperture will always be centered. In that case, the outgoing beam direction is the same as the ingoing beam direction according to Eq. (LABEL:SIEqKOutCompleteClipping), which can be rationalized because the direction of a “center beam” through a lens is unaffected by any focal length.
Finally, we need to consider the power transmission through a circular aperture. In Section LABEL:SISecFiniteAperture of the SI, we assume without loss of generality that the beam is displaced along a GOE-local $\hat{\mathbf{x}}$ coordinate with respect to the center of the partially closed GOE aperture. Then we have to solve Eq. (LABEL:SIEqPowerTransmission),
$P_{\text{out}}=I_{0}\underset{(x^{2}+y^{2}\leq a^{2})}{\int\mathrm{d}x\int\mathrm{d}y}\,\exp\left[-2\frac{(x-d)^{2}}{w^{2}}-2\frac{y^{2}}{w^{2}}\right],$ (4)
where the output power $P_{\text{out}}$ is obtained from the on-axis incident beam intensity $I_{0}$, the incident beam radius $w$, the displacement $d$, and the aperture radius $a$ through integration of the spatial Gaussian beam intensity distribution over the circular opening, which is performed numerically with pre-calculated loop-up tables to speed up the computation. In the case of a straight edge instead of a circular aperture, the power transmission is obtained in closed form in Eq. (LABEL:SIEqTransmissionStraightEdge) via the complementary error function in Eq. (LABEL:SIEqComplementaryErrorFunction).
3 N-beam interference